# Heterogeneous Treatment Effects and

### v1

@@ 6.1 Three essential theoretical questions
      For which mechanism ($j$), $X_{k}$ is a candidate MIV?
@@
      Is Assumption 2 plausible for each of the other candidate mechanisms?
@@
Question #1 is fairly standard in applied empirical research. Researchers often posit one or more mechanisms of interest in addition to alternative explanations. Questions #2 and #3, instead, are much less standard. When researchers assess HTEs, it is rare to discuss the relationship between the moderator of interest and other mechanisms even though these assumptions are necessary (but not sufficient) for mechanism attribution, as we show. Explicit justification of a moderator of interest as a candidate MIV of a mechanism may facilitate the search for MIVs. As we show in Figure 2, a MIV can moderate the effect of treatment on the mediator or the effect of the mediator on the outcome. Since the latter possibility is generally ignored in applied research, these considerations may broaden the set of candidate MIVs.
@@ 6.2 Improving the interpretation of HTEs as mechanism tests
This framework provides guidance for the interpretation of HTEs when researchers are trying to make inferences about which causal mechanisms are active. First, while the presence of HTEs provide evidence that a mechanism is active when the exclusion assumptions hold and an outcome is directly affected, the absence of HTEs is less informative (even under the same exclusion assumptions). In this sense, a lack of HTEs cannot be used to “rule out” a candidate mechanism or show that it is inert. As Proposition 2 shows, even when an outcome is directly-affected by a mechanism, a lack of HTEs could mean that (1) our model of the relationship between $X_{k}$ and mechanism $j$ is wrong; or (2) that no $MIV$ exists because the mechanism is inert or produces a homogeneous effect. Because we cannot rule explanation (1), we cannot affirm explanation (2).
This observation is particularly stark when we consider the statistical properties of HTEs. Low statistical power for interactions reduces our ability to statistically detect HTEs that do exist. In other words, we risk many false-negatives in inferences related to the existence of heterogeneity. Because a lack of HTEs is uninformative about mechanisms, low power suggests that applied researchers often operate in a world in which heterogeneity analysis is unlikely to provide information to support inferences about mechanisms. [^19] Selective reporting of significant results (due to $p$-hacking or publication bias) complicates the situation further. In this case, evidence in favor of treatment-effect heterogeneity is more likely to be a false-positive, which increases the the probability that researchers infer that a mechanism is active when it is not.
@@ 6.3 Guidance for research design
Our framework posits several recommendations for the design of causal research that seeks to test mechanisms quantitatively using HTE. Our suggestions are premised on improvements in measurement. In terms of covariates, we are primarily concerned with which covariates are measured and the number of candidate MIVs (per mechanism) among those covariates. Covariates are only useful for ascertaining mechanisms when (1) they are plausibly MIVs for a single mechanism; and (2) they do not moderate direct effects. This observation suggests that special care must be taken when positing candidate MIVs. When pre-treatment covariates are (largely) collected in baseline data collection, there is a need to posit MIVs and defend exclusion assumptions ex-ante. Such considerations require more theory and justification than are typically conveyed in the specification of moderation analyses in pre-analysis plans.
@@
When considering directly-affected outcomes, it is very useful to have multiple candidate MIVs for a given mechanism. To see why, consider the case in which we have two candidate MIVs, $X_{1}$ and $X_{2}$ for mechanism $j$, and both exclusion assumptions hold for both candidate MIVs, and the outcome is directly-affected. Suppose that there do not exist HTEs in $X_{1}$ but there do exist HTEs in $X_{2}$. If we only measured HTEs with respect to $X_{k}$, following Proposition 2, we would not be able to ascertain whether the problem is with the theory ($X_{1}\notin\textbf{X}_{j}^{MIV}$) or whether there simply exist no MIV for mechanism $j$. If there exist HTEs in $X_{2}$, we can eliminate the possibility that there do not exist MIV for mechanism $j$. This would suggest that the theory with respect to $X_{k}$ is misspecified. This is useful insofar as it allows us to make an inference that mechanism $j$ is active. Note, however, that in order to leverage multiple candidate MIVs, both exclusion assumptions must hold for each candidate MIV, which can be quite demanding.
@@
Our distinction between directly- and indirectly-affected outcomes yields two further recommendations for research design. First, if a goal of a research design is to distinguish causal mechanisms, directly-affected outcomes should be prioritized in HTE analysis. This requires researchers to make clear which outcomes are directly-affected and emphasizes the value of measuring these outcomes. In our motivating example, researchers would typically measure (self-reported) vote choice in a survey to measure the effects of a shock on incumbent support. But if they were running a survey, they could, in principle, elicit willingness-to-pay for the incumbent to try to directly measure voter utility. Our results suggest that the latter would be a worthwhile—if non-standard—investment because HTEs can provide some information about mechanisms with this latter outcome (but not the former).
@@
Second, these results merit broader consideration of latent variable measurement models (Fariss, Kenwick, and Reuning, 2020) in the case when a theorized directly-affected outcome is latent. It is rare for researchers to explicitly measure treatment effects on estimates of a latent variable (e.g., attitudes or preferences). However, popular methods for indexing multiple outcome measures, including $Z$-score indices (i.e., Kling, Liebman, and Katz, 2007), do this implicitly. More explicit consideration of which latent variables are directly affected by (a) mechanism(s) and how will improve the use of HTEs for mechanism detection. Such considerations can also provide information about which latent variable models are most appropriate.
@@ 6.4 Strengthening assumptions to use HTE as mechanism tests for indirectly-affected outcomes
First, in Appendix F.1, we show that the assumption of montonicity is not, in general, sufficient to provide information about mechanism activation through analysis of HTEs. This occurs because an assumption of montonicity fails to provide enough information about the data generating process that generates directly- and indirectly-affected outcomes. We show this in the context of the following oft-used data generating process for a binary outcome, $h(Y)\in\{0,1\}$:
@@
for some constant $c\in(-\infty,\infty)$ and random variable $\varepsilon$ distributed according to the density function $f_{\varepsilon}(\cdot)$. Under the assumption that $g$ is continuous and differentiable, this data-generating process holds that $X_{k}$ is a MIV if $\frac{\partial^{2}g}{\partial Z\partial X_{k}}\neq 0$ for some $x\in X_{k}$. An assumption of monotonicity strengthens this condition by further implying that $\frac{\partial^{2}g}{\partial Z\partial X_{k}}$ is weakly positive or negative for all $x\in X_{k}$. Our goal is to learn whether $X_{k}$ is a MIV by assessing whether $\frac{\partial^{2}g}{\partial Z\partial X_{k}}$ is non-zero through estimation of HTE.
In Proposition 6, we show that montonicity alone is not sufficient to ensure that HTEs take different signs when $X_{k}$ is not a MIV. Moreover, it does not ensure that $\frac{\partial^{2}Y}{\partial Z\partial X_{k}}>(<)0$ when $X_{k}$ is a MIV and montonicity is satisfied. These results show that we need additional assumptions on the distribution $f_{\varepsilon}$ and/or the functional form of $g(Z,X_{k})$ for monotonicity to provide sufficient information to distinguish mechanisms.
Second, and in contrast, Appendix F.2 shows how widely-used random utility models do permit inferences about mechanism activation when observed outcomes are indirectly-affected. These models provide a functional mapping between an individual’s utility and their choice between relevant alternatives. For example, in our motivating example, an individual voter’s utility from a given candidate is not generally observed whereas their self-reported vote choice is observed. A random utility model decomposes utility from a given choice into an observed systematic component and a random component. By specifying the systematic component as a function of individual- or choice-specific covariates and assuming that the random component is distributed according to a extreme value distribution, researchers can estimate the systematic component of utility. Using these estimates of our directly-affected outcome (utility), researchers can then use HTE to assess mechanism activation under Proposition 1. The invocation of a random-utility model is not free: it makes strong assumptions about the mapping between utility and choice outcomes that researchers may not be well-positioned to make or assess. However, these assumptions about the mapping between directly- and indirectly-affected outcomes do buy additional information which may permit learning about mechanism activation from HTE.
@@ 7 Conclusion
Social scientists routinely estimate HTEs with respect to a pre-treatment covariate with the stated intent of mechanism detection. We show that learning about the the activation of a mechanism using HTEs is less straightforward than conveyed by current practice. Specifically, any link between a covariate (moderator) and a mechanism requires exclusion assumptions, so that covariate does not moderate the indirect effect of other mechanisms or the direct effect of treatment on an outcome. Even when these assumptions hold, we can only use HTEs to affirm the activation of a mechanism when (1) HTEs exist and (2) the outcome is directly affected by a mechanism. Outside this case, HTEs do not provide sufficient information to show that a mechanism is active or inactive. In this sense, HTEs analysis should not be used to “rule out” mechanisms (or show that they are inert).
While mechanism detection is presently the modal use of HTEs in recent work in political science (see Table 1), it is not the only use of HTE. Our results speak to contexts where mechanistic analysis is the current aim. In current practice, HTEs are also increasingly used for extrapolation of treatment effects to different populations/settings (Egami and Hartman, 2022; Devaux and Egami, 2022) and the targeting of treatments (Athey, Tibshirani, and Wager, 2019; Kitagawa and Tetenov, 2018). Our workd oes not undermine these applications of HTEs, because these applications do not rely on questions of how to attribute observed effects to mechanisms (Slough and Tyson, 2023a).
Our analysis raises a number of issues and opportunities for future research to build upon. In particular, we emphasize the need to distinguish between the level at which mechanisms operate (e.g., on utility) and the outcomes we observe. This distinction has underappreciated implications for multiple quantitative methods to detect mechanism activation, including mediation analysis and investigation of the sign of treatment effects. In addition, our framework can help to clarify the relationship between causal mechanisms and other applications of HTEs to clarify the (applied) theoretical foundations of these approaches.
@@ Appendix A Additional Classification of Articles
Table A.1 provides an additional classification of the articles described in Table 1, by research design or identification strategy. Note that we collapse difference-in-difference and panel analyses into one category that includes two-way fixed-effects and other estimators of the average treatment effect on the treated (ATT). We also collapse IV and natural experimental analyses into a single category that includes studies with some claim of exogenous variation not created by researchers that is argued to facilitate identification of an average treatment effect (ATE); an intent-to-treat effect (ITT); or a local average treatment effect (LATE). [^20] This LATE is often termed the complier average causal effect (CACE) in political science. This table shows that using HTEs to detect mechanisms is not unique to any one research design in common usage; the proportions of articles in these journals that uses HTEs as a mechanism test (given by the “weighted average” column) is quite similar across all of these designs.
Total Pr(Reports HTE as mechanism test) Weighted
@@ B.1 Incorrect DAG
Figure A.1 depicts the DAG that is evaluated by the HTE analysis in Remarks 1-2. Note that the dashed lines do not correspond to the theoretical model. In the model, only the learning mechanism is active. This mechanism is evaluated by examining heterogeneity in treatment effects with respect to voters’ prior beliefs.
We note that this graph does not directly correspond to the model in the paper. Here, the valence shock, $v_{i}$ is measured pre-treatment (ex-ante), and the researcher (wrongly) believes that it moderates treatment effects. We represent this in the graph with $\widetilde{v}_{i}$, a measure of “ex-post” valence.
@@ B.2 Proofs of Remarks 1-2
Appendix C Directly/Indirectly-Affected versus Latent/Observed Outcomes
We propose a distinction between directly and indirectly-affected outcomes. This distinction is a theoretical claim about the causal process that generates outcomes of interest. This is different from the distinction between latent and observed outcomes. The distinction between latent and observed outcomes invokes both a theoretical claim about the causal process linking latent to observed outcomes as well as an empirical claim about the observability of the latent outcome. To show that these are distinct classifications, we propose three hypothetical studies united by a common mechanism—learning—in different contexts and organize outcomes into Table A.2. Note that we index outcomes by their sequence in the subscript and across the hypothetical studies $S1,S2,$ or $S3$ in the superscript.

### v2

@@ 6.1 Three essential theoretical questions
      For which mechanism, $M$, is $X_{k}$ a candidate MDV?
@@
      Is Assumption 2 plausible?
@@
Question #1 is fairly standard in applied empirical research. Researchers often posit one or more mechanisms of interest in addition to alternative explanations. In the exposition of our framework, we focus on one mechanism and relegate alternative explanations to the composite direct effect. C provides guidance for cases with multiple mechanisms of interest.
Questions #2 and #3, instead, are much less standard. When researchers assess HTEs, it is rare to discuss the relationship between the moderator of interest and other mechanisms (within the direct effect) even though this assumption is necessary (but not sufficient) for mechanism attribution with HTEs, as we show. Explicit justification of a moderator of interest as a candidate MDV of a single mechanism may facilitate the search for MDVs. As we show in Figure 2, a MDV can moderate the effect of treatment on the mediator or the effect of the mediator on the outcome. Since the latter possibility is generally ignored in applied research, these considerations may broaden the set of candidate MDVs.
@@ 6.2 Improving the interpretation of HTEs as mechanism tests
This framework provides guidance for the interpretation of HTEs when researchers are trying to make inferences about which causal mechanisms are active. First, while the presence of HTEs provide evidence that a mechanism is active when the exclusion assumption holds and an outcome is directly affected, the absence of HTEs is less informative (even under the same exclusion assumptions). In this sense, a lack of HTEs cannot be used to “rule out” a candidate mechanism or show that it is inert. As Proposition 2 shows, even when an outcome is directly-affected by a mechanism, a lack of HTEs could mean that (1) our model of the relationship between $X_{k}$ and mechanism $M$ is wrong; or (2) that no $MDV$ exists because the mechanism is inert or produces a homogeneous effect. Because we cannot rule explanation (1), we cannot affirm explanation (2).
This observation is particularly stark when we consider the statistical properties of HTEs. Low statistical power for interactions reduces our ability to statistically detect HTEs that do exist. In other words, we risk many false-negatives in inferences related to the existence of heterogeneity. Because a lack of HTEs is uninformative about mechanisms, low power suggests that applied researchers often operate in a world in which heterogeneity analysis is unlikely to provide information to support inferences about mechanisms. [^15] Selective reporting of significant results complicates the situation further. In this case, evidence in favor of treatment-effect heterogeneity is more likely to be a false-positive, which increases the the probability that researchers infer that a mechanism is active when it is not.
@@ 6.3 Guidance for research design
Our framework posits several recommendations for the design of causal research that seeks to test mechanisms quantitatively using HTE. Our suggestions are premised on improvements in measurement. In terms of covariates, we are primarily concerned with which covariates are measured and the number of candidate MDVs (per mechanism) among those covariates. Covariates are only useful for ascertaining mechanisms when (1) they are plausibly MDVs for a single mechanism; and (2) they do not moderate direct effects. This observation suggests that special care must be taken when positing candidate MDVs. When pre-treatment covariates are (largely) collected in baseline data collection, there is a need to posit MDVs and defend exclusion assumptions ex-ante. Such considerations require more theory and justification than are typically conveyed in the specification of moderation analyses in pre-analysis plans.
@@
When considering directly-affected outcomes, it is very useful to have multiple candidate MDVs for a given mechanism. To see why, consider the case in which we have two candidate MDVs, $X_{1}$ and $X_{2}$ for mechanism $M$, the exclusion assumption holds for both candidate MDVs, and the outcome is directly-affected. Suppose that there do not exist HTEs in $X_{1}$ but there do exist HTEs in $X_{2}$. If we only measured HTEs with respect to $X_{k}$, following Proposition 2, we would not be able to ascertain whether the problem is with the theory ($X_{1}\notin\textbf{X}^{MDV}$) or whether there simply exist no MDV for mechanism $M$. If there exist HTEs in $X_{2}$, we can eliminate the possibility that there do not exist MDV for mechanism $M$. This would suggest that the theory with respect to $X_{k}$ is misspecified. This is useful insofar as it allows us to make an inference that mechanism $M$ is active. Note, however, that in order to leverage multiple candidate MDVs, the exclusion assumption must hold for each candidate MDV, which can be quite demanding.
@@
Our distinction between directly- and indirectly-affected outcomes yields two further recommendations for research design. First, if a goal of a research design is to distinguish causal mechanisms, directly-affected outcomes should be prioritized in HTE analysis. This requires researchers to make clear which outcomes are directly-affected and emphasizes the value of measuring these outcomes. In empirical studies analogous to our motivating example, researchers would typically measure (self-reported) vote choice in a survey to measure the effects of a shock on incumbent support. But if they were running a survey, they could, in principle, elicit willingness-to-pay for the incumbent to try to directly measure voter utility. Our results suggest that the latter would be a worthwhile—if non-standard—investment because HTEs can provide some information about mechanisms with this latter outcome (but not the former).
@@
Second, these results merit broader consideration of latent variable measurement models (Fariss, Kenwick, and Reuning, 2020) in the case when a theorized directly-affected outcome is latent. It is rare for researchers to explicitly measure treatment effects on estimates of a latent variable (e.g., attitudes or preferences). However, popular methods for indexing multiple outcome measures, including $Z$-score indices (i.e., Kling, Liebman, and Katz, 2007), do this implicitly. More explicit consideration of which latent variables are directly affected by (a) mechanism(s) and how may improve the use of HTEs for mechanism detection. Such considerations can also provide information about which latent variable models are most appropriate.
@@ 6.4 Strengthening assumptions to use HTE as mechanism tests for indirectly-affected outcomes
First, in Appendix G.1, we show that the assumption of montonicity is not, in general, sufficient to provide information about mechanism activation through analysis of HTEs. This occurs because an assumption of montonicity fails to provide enough information about the data generating process that generates directly- and indirectly-affected outcomes. We show this in the context of the following oft-used data generating process for a binary outcome, $h(Y)\in\{0,1\}$:
@@
for some constant $c\in(-\infty,\infty)$ and random variable $\varepsilon$ distributed according to the density function $f_{\varepsilon}(\cdot)$. Under the assumption that $g$ is continuous and differentiable, this data-generating process holds that $X_{k}$ is a MDV if $\frac{\partial^{2}g}{\partial Z\partial X_{k}}\neq 0$ for some $x\in X_{k}$. An assumption of monotonicity strengthens this condition by further implying that $\frac{\partial^{2}g}{\partial Z\partial X_{k}}$ is weakly positive or negative for all $x\in X_{k}$. Our goal is to learn whether $X_{k}$ is a MDV by assessing whether $\frac{\partial^{2}g}{\partial Z\partial X_{k}}$ is non-zero by estimating HTEs.
In Proposition 10, we show that montonicity alone is not sufficient to ensure that HTEs take different signs when $X_{k}$ is not a MDV. Moreover, it does not ensure that $\frac{\partial^{2}Y}{\partial Z\partial X_{k}}>(<)0$ when $X_{k}$ is a MDV and montonicity is satisfied. These results show that we need additional assumptions on the distribution $f_{\varepsilon}$ and/or the functional form of $g(Z,X_{k})$ for monotonicity to provide sufficient information to distinguish mechanisms.
Second, and in contrast, Appendix G.2 shows how widely-used random utility models do permit inferences about mechanism activation when observed outcomes are indirectly-affected. These models provide a functional mapping between an individual’s utility and their choice between relevant alternatives. For example, in our motivating example, an individual voter’s utility from a given candidate is not generally observed whereas their self-reported vote choice is observed. A random utility model decomposes utility from a given choice into an observed systematic component and a random component. By specifying the systematic component as a function of individual- or choice-specific covariates and assuming that the random component is distributed according to a extreme value distribution, researchers can estimate the systematic component of utility. Using these estimates of our directly-affected outcome (utility), researchers can then use HTE to assess mechanism activation under Proposition 1. The invocation of a random-utility model is not free: it makes strong assumptions about the mapping between utility and choice outcomes that researchers may not be well-positioned to make or assess. However, these assumptions about the mapping between directly- and indirectly-affected outcomes do buy additional information which may permit learning about mechanism activation from HTE.
@@ 7 Conclusion
Social scientists routinely estimate HTEs with respect to a pre-treatment covariate with the stated intent of mechanism detection. We show that learning about the the activation of a mechanism using HTEs is far less straightforward than conveyed by current practice. Specifically, any link between a covariate (moderator) and a mechanism requires an exclusion assumption, so that covariate does not moderate other effects of treatment on an outcome. Even when this assumption holds, we can only use HTEs to affirm the activation of a mechanism when (1) HTEs exist and (2) the outcome is directly affected by a mechanism. Outside this case, HTEs do not provide sufficient information to show that a mechanism is active or inactive. In this sense, HTEs analysis should not be used to “rule out” mechanisms (or show that they are inert).
While mechanism detection is presently the modal use of HTEs in recent work in political science (see Table 1), it is not the only use of HTE. Our results speak to contexts where mechanistic analysis is the current aim. In current practice, HTEs are also increasingly used for extrapolation of treatment effects to different populations/settings (Egami and Hartman, 2022; Devaux and Egami, 2022) and the targeting of treatments (Athey, Tibshirani, and Wager, 2019; Kitagawa and Tetenov, 2018). Our work does not undermine these applications of HTEs, because these applications do not rely on questions of how to attribute observed effects to mechanisms (Slough and Tyson, 2024a).
Our analysis raises a number of issues and opportunities for future research to build upon. In particular, we emphasize the need to distinguish between the level at which mechanisms operate (e.g., on utility) and the outcomes we observe. This distinction has underappreciated implications for multiple quantitative methods to detect mechanism activation, including mediation analysis and investigation of the sign of treatment effects. Our framework can help to clarify the relationship between causal mechanisms and other applications of HTEs to clarify the (applied) theoretical foundations of these approaches.
@@ Appendix A Additional Classification of Articles
Table A.1 provides an additional classification of the articles described in Table 1, by research design. Note that we collapse difference-in-difference and panel analyses into one category that includes two-way fixed-effects and other estimators of the average treatment effect on the treated (ATT). We also collapse IV and natural experimental analyses into a single category that includes studies with some claim of exogenous variation not created by researchers that is argued to facilitate identification of an average treatment effect (ATE); an intent-to-treat effect (ITT); or a local average treatment effect (LATE). [^16] This LATE is often termed the complier average causal effect (CACE) in political science. This table shows that using HTEs to detect mechanisms is not unique to any one research design in common usage; the proportions of articles in these journals that uses HTEs as a mechanism test (given by the “weighted average” column) is quite similar across all of these designs.
Pr(Reports HTEs as
Total mechanism test) Weighted
@@ B.1 Incorrect DAG
Figure A.1 depicts the DAG that is evaluated by the HTE analysis in Remarks 1-2. Note that the dashed lines are absent in the theoretical model. In the model, only the learning mechanism is active. This mechanism is evaluated by examining heterogeneity in treatment effects with respect to voters’ prior beliefs.
We note that this graph does not precisely correspond to the model in the paper. Here, the valence shock, $v_{i}$ is measured pre-treatment (ex-ante), and the researcher (wrongly) believes that it moderates treatment effects. We represent this in the graph with $\widetilde{v}_{i}$, a measure of “ex-post” valence.
@@ B.2 Proofs of Remarks 1-2
Appendix C Generalized Framework with Multiple Mechanisms
Suppose that there exist $J\geq 1$ mediators (or mechanisms), indexed by $M_{1},...,M_{J}$. Given two treatment values, $z,z^{\prime}\in Z$, the total effect of $Z$ on $Y$ is:
Our notation varies slightly from conventional presentations of mediation that only consider two mechanisms (mediators) (Imai, Keele, and Tingley, 2010). To this end, $j-$ and $j+$ denote index $h\in J$ such that $h<j$ and $h>j$ respectively. Treatment effects may consist of direct ($DE$) and indirect ($IE_{j}$) effects, as follows: [^17] The intuition for our notation is as follows: define $IE_{0}(z,z^{\prime};X)=DE(z,z^{\prime};X)$. This implies that the first term of $IE_{j}(z,z^{\prime};X)$ and the second term of $IE_{j-1}(z,z^{\prime};X)$ cancel out.
$Y(z^{\prime},M_{j-}(z^{\prime};X),M_{j}(z^{\prime};X),M_{j+}(z;X);X)$
The direct effect, $DE(z,z^{\prime};X)$ represents the direct effect of $Z$ on $Y$ holding mediators at potential outcomes $M_{j}(z;X)$. It is not necessary to believe that treatments produce unmediated (direct) effects on outcomes to use this framework. We allow for direct effects in the interest of generality. The indirect effects, measures the effect on the outcome that operates by changing the potential outcome of the mediator. As is standard, we can then re-write the total effect as follows:
which is defined at the unit, or individual level. Although there exists multiple ways to decompose total effect and this decomposition holds mechanisms to be independent, all results in the paper hold if other mediators (other than mediator $j$ of interest) are arbitrarily correlated.
If we evaluate expectations over $X$, we obtain:
$ATE(z,z^{\prime})$ $=E_{X}[Y(z)-Y(z^{\prime})]$ (32)
$=E_{X}[DE(z,z^{\prime};X)+\sum_{j=1}^{J}IE_{j}(z,z^{\prime};X)]$ (33)
$=ADE(z,z^{\prime})+\sum_{j=1}^{J}AIE_{j}(z,z^{\prime})$ (34)
We use $ADE$ and $AIE$ to denote average direct effect and average indirect effect.
Because there are multiple mediators, we require a further assumption beyond 2 (reproduced from the main text below). Assumption 3 rules out the possibility that the indirect effects produced by other mechanisms are moderated by $X_{k}$.
Assumption 2 (Exclusion I).
Assumption 3 (Exclusion II).
Assumptions 2-3 constrain the relationship between a moderator, $X_{k}$, and other mechanisms/the direct effect. Figure A.2 depicts violations of Assumptions 2-3 graphically for a MDV ($X^{MDV}$) of mechanism 1, $M_{1}$. Assumption 2 rules out the blue dashed path from $X^{MDV}$ to the direct effect of $Z$ on $Y$. Assumption 3 rules out either/both red dot-dashed path from $X^{MDV}$ to the effect of $Z$ on a different mechanism, $M_{2}$. This means that $X_{k}$ does not moderate the effect of $Z$ on $M_{2}$. Nor does $X_{k}$ moderate the effect of $M_{2}$ on $Y$. However, there can be a direct arrow from $X_{k}$ to $M_{2}$. This means that $X_{k}$ can cause (or predict) the level of $M_{2}$, but it cannot interact with treatment in any way to cause $M_{2}$ or the effect of $M_{2}$ on $Y$. These assumptions are crucial for the use of HTEs to detect mechanisms. In their absence, we cannot link a heterogeneous treatment effect to a unique theoretical mechanism representation (mediator) because we cannot ensure additive separability of the effect of the mechanism of interest from the effects of other (possible) mechanisms.
[…]
Figure A.2: Assumption 2 rules out the blue dashed path. Assumption 3 rules out both of the red dot-dashed paths. All orange and black solid paths are permissible under Assumptions 2 and 3. For $X_{k}$ to be an MDV, at least one of the two orange paths must be present.
Appendix D Directly/Indirectly-Affected versus Latent/Observed Outcomes
We propose a distinction between directly and indirectly-affected outcomes. This distinction is a theoretical claim about the causal process that generates outcomes of interest. This is different from the distinction between latent and observed outcomes. The distinction between latent and observed outcomes invokes both a theoretical claim about the causal process linking latent to observed outcomes as well as an empirical claim about the (non)-observability of the latent outcome. To show that these are distinct classifications, we propose three hypothetical studies united by a common mechanism—learning—in different contexts and organize outcomes into Table A.2. Note that we index outcomes by their sequence in the subscript and across the hypothetical studies $S1,S2,$ or $S3$ in the superscript.